# Generalizing matrix representations to fully heterochronous ranked tree shapes

### v3

@@ Abstract
Phylogenetic tree shapes capture fundamental signatures of evolution. We consider “ranked” tree shapes, which are equipped with a total order on the internal nodes compatible with the tree graph. Recent work has established an elegant bijection of ranked tree shapes and a class of integer matrices, called F-matrices, defined by simple inequalities. This formulation is for isochronous ranked tree shapes, where all leaves share the same sampling time, such as in the study of ancient human demography from present-day individuals. Another important style of phylogenetics concerns trees where the “timing” of events is by branch length rather than calendar time. This style of tree, called a rooted phylogram, is output by popular maximum-likelihood methods. These trees are broadly relevant, such as to study the affinity maturation of B cells in the immune system. Discretizing time in a rooted phylogram gives a fully heterochronous ranked tree shape, where leaves are part of the total order. Here we extend the F-matrix framework to such fully heterochronous ranked tree shapes. We establish an explicit bijection between a class of F-matrices and the space of such tree shapes. The matrix representation has the key feature that values at any entry are highly constrained via four previous entries, enabling straightforward enumeration of all valid tree shapes. We also use this framework to develop probabilistic models on ranked tree shapes. Our work extends understanding of combinatorial objects that have a rich history in the literature.
@@ Introduction
Evolution is the unifying theme of biology, and it operates in diverse modes. These modes can be seen in the structure of phylogenetic trees (MooersHeard). For example, the tree of influenza has a highly “imbalanced” shape, which comes from intense evolutionary selective pressure from host immunity, in contrast with the trees of other viruses Grenfell2004-dz. Scientists characterize these modes of evolution by studying phylogenetic tree “shapes”: rooted bifurcating tree graphs without leaf labels.
@@
An elegant means of characterizing tree shapes has recently been developed, which includes information about relative ordering of nodes in addition to graph structure Kim2020-ip; Samyak2024-yo. This relative ordering is expressed in terms of a ranking, i.e., a total ordering of the internal nodes of the tree. The combination of the tree shape and relative ordering defines a “ranked tree shape.” There is a bijection between such ranked tree shapes and a class of integer-valued matrices, called “F-matrices”, which are characterized by simply-expressed inequalities Kim2020-ip; Samyak2024-yo. By recording information about the order of events on the tree, this formulation enables richer comparison than tree structure alone. However, the existing formulation of F-matrices is limited to “isochronous” ranked tree shapes (Figure 1, left) in which all the leaves of the tree are assumed to have been sampled at the same time, or at least at known fixed sampling times. This makes perfect sense in the setting of “time trees” (a.k.a. chronograms): phylogenetic trees with nodes labeled by calendar time and leaf nodes representing molecular sequences with known sampling times. Such trees result from inference done using software such as BEAST Drummond2007-co; Bouckaert2019-wv; Baele2025-fe or TreeTime Sagulenko2018-xl.
@@
Figure 1: Left: an isochronous tree shape; an F-matrix bijection has been established for such objects Kim2020-ip; Samyak2024-yo. Right: a fully heterochronous tree shape; the present manuscript establishes an analogous bijection between these objects and a class of F-matrices. The two trees are isomorphic as graphs, but are not the same type of ranked tree shape. On the left isochronous tree, internal nodes have unique ranks and leaves share a common rank. We mark leaves in gray to indicate that they do not form part of the “data” encoded by the ranked tree. On the right fully heterochronous tree, all nodes have unique ranks and the rank of a leaf may be less than the rank of an internal node.
@@
There is another type of tree analysis that simply represents the phylogenetic tree without timing constraints, letting the length of each edge represent the amount of evolution that has happened along that edge. This structure is called a “rooted phylogram”. Phylograms are the inferential output of software such as IQ-TREE Minh2020-va and RAxML Stamatakis2014-xo.
@@
One may wish to use rooted phylograms to study patterns of evolution in systems where dates are not available or relevant. For example, in B cell affinity maturation, the evolutionary structure of the phylogenetic tree is determined by a relatively short period in the germinal center, after which the resulting cells circulate for longer as memory B cells without further mutation Victora2022-fm. Due to this two-part process the blood sampling time for B cells is not relevant to the actual “sampling time” of the B cells, which is the various times when they left the germinal center. Hence, one can use a rooted phylogram.
@@
Crucially, the leaf positions of rooted phylogram inference form part of the inferential output, in contrast to time tree inference (for which they form part of the input data). Thus, we wish to capture the positions of the leaf nodes as part of our tree representation. As with the time tree case, we discretize the positions of the internal nodes into a ranking, obtaining what we call a fully heterochronous ranked tree shape (Figure 1, right).
@@
In this paper, we extend the previous F-matrix characterization to fully heterochronous ranked tree shapes and prove theorems on matrix construction. Going further, we provide a method to iteratively build all F-matrices one entry at a time. This extends previous work in two ways. First, in (Kim2020-ip), the authors proposed a bijective F-matrix encoding of ranked tree shapes (isochronous and heterochronous) to define a distance on the space via matrix norms. This defined F-matrices by the ranked tree shapes that they encode, while the set of matrices comprising F-matrices was explicitly identified for isochronous ranked tree shapes only (Samyak2024-yo); we restate the latter result as Theorem 1. Second, an iterative construction was noted for the isochronous case, but it was not explicitly stated nor was its correctness proved. In both the isochronous and heterochronous case, F-matrices are classified by their entries satisfying a number of simple linear inequalities. The iterative construction is a method to solve the linear inequalities without the need for back-substitution. This construction also yields an explicit enumeration of all F-matrices, and as such all fully heterochronous ranked tree shapes.
@@
The current literature lacks descriptive probability distributions on the space of fully heterochronous ranked tree shapes. To address this, we first introduce two parameter-free models: a backward-in-time coalescent model (kingman1982coalescent), and a forward-in-time model referred to here as diagonal top-down. These may be considered as null models. In the opposite direction, we exploit the iterative matrix construction to define highly flexible probability distributions with many free parameters on the space of fully heterochronous ranked tree shapes. This general construction can be specialized to a particular class of beta-splitting model (aldous1996probability; sainudiin2016beta). Future work will focus on fitting these flexible distributions via neural networks.
@@ 1 Definitions and connection with previous literature
- Discretized inferential output of e.g. BEAST or TreeTime. - Branch lengths are in units of calendar time. - Internal nodes are totally ordered. - Leaf positions are part of the input data for inference. - Discretized inferential output of e.g. IQ-TREE or RAxML. - Branch lengths are in units of evolutionary change. - All nodes are totally ordered. - The leaf positions are part of the inferential output.
@@
Ranked tree shapes are related to another type of tree structure, which are known by many names including binary increasing trees, ordered binary trees, or André trees foata1971; donaghey1975; poupard1989. Ordered (increasing, André) binary trees are fully heterochronous ranked tree shapes without the assumption that the binary tree is full; such trees have nodes with out-degree at most two instead of out-degree exactly zero or two. Fully heterochronous ranked tree shapes are also called strictly ordered binary trees. While isochronous ranked tree shapes are not ordered binary trees, the isochronous ranked tree shapes with $n$ leaves are equinumerous with the ordered binary trees with n1n-1 nodes. Ordered binary trees are inherently related to alternating permutations that were extensively studied in andre1881, which is why some authors call such trees André trees.
Let $\mathcal{T}_{n}$ denote the set of isochronous ranked tree shapes with $n$ leaves and $\mathcal{T}^{*}_{n}$ denote the set of fully heterochronous ranked tree shapes with $n$ leaves. The cardinalities of these sets correspond to the so called Euler up/down (or zigzag) numbers and reduced tangent numbers OEIS-iso; OEIS-hetero. In particular, in terms of exponential generating functions, we have
@@ 2 Preliminaries
For a ranked tree shape we define three types of matrices, which we call F, D, and E-matrices. The differences between these matrices and their isochronous analogs are minor, and we highlight where differences occur. One additional difference with previous work is that we will use the convention that indices start at 0, not $1$, in order to make theorem statements cleaner. In formulating the matrices, we give a purely graph theoretic definition and then an interpretation where the total ordering is based on events occurring in time, which is relevant for applications and is useful when visualizing such trees.
@@
The F-matrix associated to $T$ is a lower triangular matrix $F$, where the size of the matrix is $(n-1)\times(n-1)$ in the isochronous case and $(2n-2)\times(2n-2)$ in the fully heterochronous case. The entry $F_{i,j}$, for $0\leq j\leq i$, is defined as the number of edges from nodes $v$ to nodes $w$, where the rank of $v$ is at most $j$ and the rank of $w$ is larger than $i$. The associated D and E-matrices are also lower triangular matrices and of the same size as the F-matrix. The entry $D_{i,j}$ is defined as the number of edges descending from the node with rank $j$ to nodes with rank larger than $i$. The entry $E_{i,j}$ is defined as the number of edges from the node with rank $j$ to the node(s) with rank $i+1$.
@@
While not labeled in the figure, the leaves are viewed as having rank 4. The associated matrices are:
@@
For a description of these matrices in line with their introduction in Kim2020-ip; Samyak2024-yo, we view $T$ as describing a branching and sampling process of lineages over time. As in previous literature, we take the approach that time moves in the direction of leaf to root (for instance, one might think of the time units being in millions of years ago). In the isochronous case, we take real numbers $u_{0}>u_{1}>\dotsb>u_{n-1}=0$ and say the event for the node(s) with rank $i$ occurs exactly at time $u_{i}$. In the fully heterochronous case we instead take real numbers $u_{0}>u_{1}>\dotsb>u_{2n-2}$, as there are more ranked nodes in this case. No event occurs in any time interval $(u_{i},u_{i+1})$. Exactly one event occurs at each time $u_{i}$, except for time $u_{n-1}$ in the isochronous case.
The entries of such matrices are non-negative integers. Given that the trees are binary, the entries of a D-matrix are restricted to $\{0,1,2\}$. In the fully heterochronous case, the entries of a E-matrix are restricted to $\{0,1\}$. In the isochronous case, the entries of all but the last row of a E-matrix are restricted to $\{0,1\}$, while entries of the last row are restricted to $\{0,1,2\}$ (as the leaves share a common rank).
@@
Kim2020-ip; Samyak2024-yo The space of isochronous ranked tree shapes with $n$ leaves is in bijection with the space of $(n-1)\times(n-1)$ F-matrices, which are lower triangular square matrices of nonnegative integers that obey the following constraints.
@@
A consequence of Theorem 1 is that it allows us to enumerate the whole space of isochronous ranked tree shapes, with a fixed number of leaves, via $\mathbf{F}$-matrices. The values of the diagonal and subdiagonal entries are common to all $\mathbf{F}$-matrices. The $\mathbf{F}$-matrices are enumerated by then selecting values for the remaining lower diagonal entries in lexicographical order, which is the order of rows then columns (see Example 3). As it turns out, selecting values in this order not only produces all F-matrices, but also never produces an invalid matrix. That is to say, setting $F_{i,j}$ to either $\min\left(F_{i-1,j},F_{i,j-1}+F_{i-1,j}-F_{i-1,j-1}\right)$ or $\max\left(F_{i,j-1},F_{i-1,j}-1,F_{i,j-1}+F_{i-1,j}-F_{i-1,j-1}-1\right)$ does not yield an unsatisfiable system of inequalities for entries filled after $F_{i,j}$.
@@ 3 Theorems
We note that the $\mathbf{F}$-matrix of a fully heterochronous ranked tree with $n$ leaves is a matrix of dimension 2n22n-2 with a different constraint on the diagonal from the isochronous case. Recall the $i$-th diagonal entry indicates the number of lineages (or edges) extant at the $i$-th time epoch and so the diagonal entries either increase by one or decrease by one depending on whether the $i$-th node is of out-degree 2 or of out-degree 0. In the isochronous case, diagonal entries always increase by one. In the following theorem, we classify F-matrices of fully heterochronous ranked tree shapes in terms of a system of inequalities.
@@
Condition 1 is equivalent to $D_{i,j}\geq 0$, which is true. Condition 2 states that the number of edges from nodes $v$ to nodes $w$, where $\text{rank}(v)\leq j$ and $\text{rank}(w)=i$, is exactly 1 or 0 (either the parent node of $w$ has rank at most $j$ or not).
We handle each part of condition 3 in order of appearance. Since the root node has exactly two children, $F_{0,0}=2$. The same edges are counted by $F_{i-1,i-1}$ and $F_{i,i}$ except for three: the edge to the node with rank $i$ (counted by $F_{i-1,i-1}$) and the two edges from the node with rank $i$ (counted by $F_{i,i}$, if they exist), so $F_{i,i}-F_{i-1,i-1}=\pm 1$. There is a single node with rank larger than 2n32n-3, so $F_{2n-3,2n-3}=1$. The same edges are counted by $F_{i-1,i-1}$ and $F_{i,i-1}$ except the edge to the node of rank $i$ (counted by $F_{i-1,i-1}$), so $F_{i-1,i-1}-F_{i,i-1}=1$. Condition 33(c) is equivalent to $E_{i-1,j}\in\{0,1\}$, which is true.
Next we prove that the conditions are sufficient. It is easier to work with the $\mathbf{D}$ and $\mathbf{E}$ matrices, rather than work directly with the $\mathbf{F}$-matrix. Suppose $F$ is a matrix satisfying the conditions in the statement of the theorem. Let $D$ and $E$ be the matrices defined by (2). We show that $E$ is the offset adjacency matrix of some totally ranked tree shape. This requires verifying the following conditions for $E$:
@@
We can determine all F-matrices by filling the remaining entries in order of $F_{2,0}$, $F_{3,0}$, and $F_{3,1}$. For $F_{2,0}$ we have two options, 0 or 1. Suppose we select $F_{2,0}=0$. Moving to $F_{3,0}$, we are forced to select $F_{3,0}=0$ by constraint 2. Lastly, for $F_{3,1}$ our options are 1 or 2, both of which yield valid F-matrices. One can verify that if we instead begin with $F_{2,0}=1$, the remaining entries work out in a similar fashion.
@@
For $F_{2,0}$ we have two options, 0 or 1, suppose we take $F_{2,0}=0$. We are forced to have $F_{2,1}=2$ by constraint 33(b). Next we must take $F_{2,2}=2$, as $F_{2,2}=4$ yields the contradiction $3=F_{3,2}\leq F_{3,3}=1$ by constraints 3(a,b). Additionally, we are forced to take $F_{3,0}=0$ by constraint 2. In
@@
This example shows how the strategy of filling rows in order of top to bottom and left to right produces all F-matrices, in both the isochronous and heterochronous case, but additional constraints are necessary to prevent invalid F-matrices in the heterochronous case. Specifically, some combinations of values for $F_{i,j}$ and $F_{i,i-1}$ from items 33(c) and 33(b) in Theorem 2 may conflict with item 1. In the example, the invalid combination is $F_{i,j}=F_{3,1}=2$ and $F_{i,i-1}=F_{3,2}=1$.
For the remainder of this section, we describe a matrix-filling strategy that does not lead to contradictions in the heterochronous case. As the subdiagonal entries are determined by the diagonal entries, we first verify that any choice of diagonal entries by item 33(a) and an additional constraint yields at least one valid $\mathbf{F}$-matrix. That is to say, when solving the system of inequalities in Theorem 2, we may select values for the diagonal without backtracking.
@@
We show how to extend an $(n,N,M)$ F-sequence to an F-matrix and that every F-matrix appears in such a way.

### v4

@@ Abstract
Phylogenetic tree shapes capture fundamental signatures of evolution. We consider “ranked” tree shapes, which are equipped with a total order on the internal nodes compatible with the tree graph. Recent work has established an elegant bijection between ranked tree shapes and a class of integer matrices, called F-matrices, defined by simple inequalities. This formulation is for isochronous ranked tree shapes, where all leaves share the same sampling time, such as in the study of ancient human demography from present-day individuals. However, branch lengths of phylogenetic trees can represent units other than calendar time, such as evolutionary distance. A tree equipped with branch lengths quantifying evolutionary distance, called a rooted phylogram, is output by popular maximum-likelihood methods. These trees are broadly relevant, such as to study the affinity maturation of B cells in the immune system. Discretizing time in a rooted phylogram gives a fully heterochronous ranked tree shape, where leaves are part of the total order. Here we extend the F-matrix framework to such fully heterochronous ranked tree shapes. We establish an explicit bijection between a class of F-matrices and the space of such tree shapes. The matrix representation has the key feature that the value at any entry is highly constrained by four previous entries, enabling straightforward enumeration of all valid tree shapes. We also use this framework to develop probabilistic models on ranked tree shapes. Our work extends understanding of combinatorial objects that have a rich history in the literature.
@@ Introduction
Evolution is the unifying theme of biology, and it operates in diverse modes. These modes can be seen in the structure of phylogenetic trees (Mooers and Heard, 1997). For example, the tree of influenza has a highly “imbalanced” shape, which comes from intense evolutionary selective pressure from host immunity, in contrast with the trees of other viruses Grenfell et al. (2004). Scientists characterize these modes of evolution by studying phylogenetic tree “shapes”: rooted bifurcating tree graphs without leaf labels.
@@
An elegant means of characterizing tree shapes has recently been developed that includes information about the relative ordering of nodes in addition to graph structure Kim et al. (2020); Samyak and Palacios (2024). This relative ordering is expressed as a ranking, i.e., a total ordering of the internal nodes of the tree. The combination of tree shape and relative ordering defines a “ranked tree shape.” There is a bijection between such ranked tree shapes and a class of integer-valued matrices, called “F-matrices”, characterized by simply-expressed inequalities Kim et al. (2020); Samyak and Palacios (2024). By recording information about the order of events on the tree, this formulation enables richer comparison than tree structure alone. However, the existing formulation of F-matrices is limited to “isochronous” ranked tree shapes (Figure 1, left) in which all the leaves of the tree are assumed to have been sampled at the same time, or at least at known fixed sampling times. This makes perfect sense in the setting of “time trees” (a.k.a. chronograms): phylogenetic trees with nodes labeled by calendar time and leaf nodes representing molecular sequences with known sampling times. Such trees result from inference done using software such as BEAST Drummond and Rambaut (2007); Bouckaert et al. (2019); Baele et al. (2025) or TreeTime Sagulenko et al. (2018).
@@
Figure 1: Left: an isochronous tree shape; an F-matrix bijection has been established for such objects Kim et al. (2020); Samyak and Palacios (2024). Right: a fully heterochronous tree shape; the present manuscript establishes an analogous bijection between these objects and a class of F-matrices. The two trees are isomorphic as graphs, but are not the same type of ranked tree shape. On the left isochronous tree, internal nodes have unique ranks and leaves share a common rank. We mark leaves in gray to indicate that they do not form part of the “data” encoded by the ranked tree. On the right fully heterochronous tree, all nodes have unique ranks and the rank of a leaf may be less than the rank of an internal node.
@@
There is another type of tree analysis that simply represents the phylogenetic tree without timing constraints, letting the length of each edge represent the amount of evolution that has happened along that edge. This structure is called a “rooted phylogram”. Phylograms are the inferential output of software such as IQ-TREE Minh et al. (2020) and RAxML Stamatakis (2014).
@@
One may wish to use rooted phylograms to study patterns of evolution in systems where dates are not available or relevant. For example, in B cell affinity maturation, the evolutionary structure of the phylogenetic tree is determined by a relatively short period in the germinal center, after which the resulting cells circulate for longer as memory B cells without further mutation Victora and Nussenzweig (2022). Due to this two-part process, the blood sampling time for B cells is not relevant to the actual “sampling time” of the B cells, which corresponds to the various times when they left the germinal center. Hence, one can use a rooted phylogram.
@@
Crucially, the leaf positions of rooted phylogram inference form part of the inferential output, in contrast to time tree inference (for which they form part of the input data). Thus, we wish to capture the positions of the leaf nodes as part of our tree representation. Consequently, we discretize the positions of all nodes into a total ranking, obtaining what we call a fully heterochronous ranked tree shape (Figure 1, right).
@@
In this paper, we extend the previous F-matrix characterization to fully heterochronous ranked tree shapes and additionally prove theorems for matrix construction. Taking this further, we provide a method to iteratively construct all F-matrices one entry at a time. This extends previous work in two ways. First, in (Kim et al., 2020), the authors proposed a bijective F-matrix encoding of ranked tree shapes (isochronous and heterochronous) to define a distance on the space via matrix norms. This defined F-matrices by the ranked tree shapes that they encode, while the set of matrices comprising F-matrices was explicitly identified for isochronous ranked tree shapes only (Samyak and Palacios, 2024); we restate the latter result as Theorem 1. Second, an iterative construction was noted for the isochronous case, but it was not explicitly stated nor was its correctness proved. As we will show, in both the isochronous and heterochronous case, F-matrices are classified by their entries satisfying a number of simple linear inequalities. The proposed iterative matrix construction is a method to solve the linear inequalities without the need for back-substitution. This construction also yields an explicit enumeration of all F-matrices, and as such all fully heterochronous ranked tree shapes. Additionally, this construction facilitates the definition of novel probabilistic models.
@@
Probability distributions on trees are important in phylogenetics, such as for maximum likelihood estimation and as priors in Bayesian inference. Examples of widely used probability distributions on isochronous phylogenetic trees include Kingman’s coalescent model on ranked labeled tree shapes and Tajima’s coalescent model on ranked unlabeled tree shape (Kingman, 1982; Tajima, 1983). The current literature lacks descriptive probability distributions on the space of fully heterochronous ranked tree shapes beyond birth-death models that assume a uniform tree topology. To address this, we first introduce two parameter-free models: a backward-in-time coalescent model (Kingman, 1982), and a forward-in-time model referred to here as diagonal top-down. These may be considered as null models. In the opposite direction, we exploit the iterative matrix construction to define highly flexible probability distributions with many free parameters on the space of fully heterochronous ranked tree shapes. This general construction can be specialized to a particular class of beta-splitting model (Aldous, 1996; Sainudiin and Véber, 2016). Future work will focus on fitting these flexible distributions via neural networks.
@@ 1 Definitions and connection with previous literature
- Discretized inferential output of e.g. BEAST or TreeTime. - Branch lengths are in units of calendar time. - Internal nodes are totally ordered. - Leaf positions are assumed to be equal and known. - Discretized inferential output of e.g. IQ-TREE or RAxML. - Branch lengths are in units of evolutionary change. - All nodes are totally ordered. - The leaf positions are part of the inferential output.
@@
Ranked tree shapes are related to another type of tree structure known by many names including binary increasing trees, ordered binary trees, or André trees Dominique Foata (1971); Donaghey (1975); Poupard (1989). Ordered (increasing, André) binary trees are fully heterochronous ranked tree shapes without the full-binary-tree requirement; such trees have nodes with out-degree at most two instead of out-degree exactly zero or two. Fully heterochronous ranked tree shapes are also called strictly ordered binary trees. While isochronous ranked tree shapes are not ordered binary trees, the isochronous ranked tree shapes with $n$ leaves are equinumerous with the ordered binary trees with n1n-1 nodes. Ordered binary trees are inherently related to alternating permutations that were extensively studied in André (1881), which is why some authors call such trees André trees. In the phylogenetics literature, Gavryushkina et al. (2013) developed efficient algorithms for counting fully heterochronous ranked tree shapes using a coalescent-based recursion, and extended this enumeration to serial sampling scenarios where samples are collected at multiple time points.
Let $\mathcal{T}_{n}$ denote the set of isochronous ranked tree shapes with $n$ leaves and $\mathcal{T}^{*}_{n}$ denote the set of fully heterochronous ranked tree shapes with $n$ leaves. The cardinalities of these sets correspond to the so-called Euler up/down (or zigzag) numbers and reduced tangent numbers OEIS Foundation Inc. (2025) ; OEIS Foundation Inc. (2025) . In particular, in terms of exponential generating functions, we have
@@ 2 Preliminaries
For a ranked tree shape we define three types of matrices, which we call F-, D-, and E-matrices. The differences between these matrices and their isochronous analogs are minor, and we highlight where differences occur. One additional difference with previous work is that we will use the convention that indices start at 0, not $1$, in order to make theorem statements cleaner. In formulating the matrices, we give a purely graph theoretic definition and then an interpretation where the total ordering is based on events occurring in time, which is relevant for applications and is useful when visualizing such trees.
@@
The F-matrix associated to $T$ is a lower triangular matrix $F$, where the size of the matrix is $(n-1)\times(n-1)$ in the isochronous case and $(2n-2)\times(2n-2)$ in the fully heterochronous case.
Definition 1.
The entry $F_{i,j}$, for $0\leq j\leq i$, is defined as the number of edges from nodes $v$ to nodes $w$, where the rank of $v$ is at most $j$ and the rank of $w$ is larger than $i$.
The associated D and E-matrices are also lower triangular matrices and of the same size as the F-matrix. The entry $D_{i,j}$ is defined as the number of edges descending from the node with rank $j$ to nodes with rank larger than $i$. The entry $E_{i,j}$ is defined as the number of edges from the node with rank $j$ to the node(s) with rank $i+1$.
For a description of these matrices in line with their introduction in Kim et al. (2020); Samyak and Palacios (2024), we view $T$ as describing a branching and sampling process of lineages over time. A coalescent or branching event corresponds to an internal node in $T$, while a sampling event corresponds to a leaf in $T$. As in previous literature, we take the approach that time moves in the direction of leaf to root (for instance, one might think of the time units being in millions of years ago). In the isochronous case, we take real numbers $u_{0}>u_{1}>\dotsb>u_{n-1}=0$ and say the event for the node(s) with rank $i$ occurs exactly at time $u_{i}$. In the fully heterochronous case we instead take real numbers $u_{0}>u_{1}>\dotsb>u_{2n-2}$, as there are more ranked nodes in this case. No event occurs in any time interval $(u_{i},u_{i+1})$. Exactly one event occurs at each time $u_{i}$, except for time $u_{n-1}$ in the isochronous case.
@@
Let us examine the second column of these matrices. For the F-matrix, we consider the three lineages present just before time $u_{1}$ (i.e., those below the dashed line labeled $u_{1}$). Of these three lineages, all are present by time $u_{2}$, so $F_{1,1}=3$; two are present by time $u_{3}$, so $F_{2,1}=2$; and one is present by time $u_{4}$, so $F_{3,1}=1$. For the D-matrix, we consider the two lineages that appear at time $u_{1}$ (i.e., the lineages with ranks $2$ and $4$). Of these two lineages, both are present by time $u_{2}$, so $D_{1,1}=2$; one is present by time $u_{3}$, so $D_{2,1}=1$; and one is present by time $u_{4}$, so $D_{3,1}=1$. For the E-matrix, we again consider the two lineages that appear at time $u_{1}$. Of these two lineages, one is sampled at time $u_{2}$, so $E_{1,1}=1$; none are sampled at time $u_{3}$, so $E_{2,1}=0$; and one is sampled at time $u_{4}$, so $E_{3,1}=1$.
@@
Note that all the leaves have rank 4. The associated matrices are:
@@
The entries of such matrices are non-negative integers. Given that the trees are binary, the entries of a D-matrix are restricted to $\{0,1,2\}$. In the fully heterochronous case, the entries of an E-matrix are restricted to $\{0,1\}$. In the isochronous case, the entries of all but the last row of an E-matrix are restricted to $\{0,1\}$, while entries of the last row are restricted to $\{0,1,2\}$ (as the leaves share a common rank).
@@
Kim et al. (2020); Samyak and Palacios (2024) The space of isochronous ranked tree shapes with $n$ leaves is in bijection with the space of $(n-1)\times(n-1)$ F-matrices, which are lower triangular square matrices of nonnegative integers that obey the following constraints.
@@
A consequence of Theorem 1 is that it allows us to enumerate the whole space of isochronous ranked tree shapes, with a fixed number of leaves, via $\mathbf{F}$-matrices. The values of the diagonal and subdiagonal entries are common to all $\mathbf{F}$-matrices. The $\mathbf{F}$-matrices are enumerated by then selecting values for the remaining lower diagonal entries in lexicographical order, which is the order of rows then columns (see Example 3). As it turns out, selecting values in this order not only produces all F-matrices, but also never produces an invalid matrix. That is to say, setting $F_{i,j}$ to either $\min\left(F_{i-1,j},F_{i,j-1}+F_{i-1,j}-F_{i-1,j-1}\right)$ or $\max\left(F_{i,j-1},F_{i-1,j}-1,F_{i,j-1}+F_{i-1,j}-F_{i-1,j-1}-1\right)$ never yields an unsatisfiable system of inequalities for entries filled after $F_{i,j}$.
@@ 3 Theorems
We note that the $\mathbf{F}$-matrix of a fully heterochronous ranked tree with $n$ leaves is a matrix of dimension $(2n-2)\times(2n-2)$ with a different constraint on the diagonal from the isochronous case. Recall the $i$-th diagonal entry indicates the number of lineages (or edges) extant at the $i$-th time epoch and so the diagonal entries either increase by one or decrease by one depending on whether the $i$-th node is of out-degree 2 or of out-degree 0. In the isochronous case, diagonal entries always increase by one. In the following theorem, we classify F-matrices of fully heterochronous ranked tree shapes in terms of a system of inequalities.
@@
Since $F_{i,j}-F_{i,j-1}=D_{i,j}$, Condition 1 is equivalent to $D_{i,j}\geq 0$, which is true. Condition 2 states that the number of edges from nodes $v$ to nodes $w$, where $\text{rank}(v)\leq j$ and $\text{rank}(w)=i$, is exactly 1 or 0 (either the parent node of $w$ has rank at most $j$ or not).
We handle each part of Condition 3 in order of appearance. Since the root node has exactly two children, $F_{0,0}=2$. The same edges are counted by $F_{i-1,i-1}$ and $F_{i,i}$ except for three: the edge to the node with rank $i$ (counted by $F_{i-1,i-1}$) and the two edges from the node with rank $i$ (counted by $F_{i,i}$, if they exist), so $F_{i,i}-F_{i-1,i-1}=\pm 1$. There is a single node with rank larger than 2n32n-3, so $F_{2n-3,2n-3}=1$. The same edges are counted by $F_{i-1,i-1}$ and $F_{i,i-1}$ except the edge to the node of rank $i$ (counted by $F_{i-1,i-1}$), so $F_{i-1,i-1}-F_{i,i-1}=1$. Since $F_{i,j-1}-F_{i,j}+F_{i-1,j}-F_{i-1,j-1}=D_{i-1,j}-D_{i,j}=E_{i-1,j}$, Condition 33(c) is equivalent to $E_{i-1,j}\in\{0,1\}$, which is true.
Next we prove that the conditions are sufficient. It is easier to work with the $\mathbf{D}$- and $\mathbf{E}$-matrices than directly with the $\mathbf{F}$-matrix. Suppose $F$ is a matrix satisfying the conditions in the statement of the theorem. Let $D$ and $E$ be the matrices defined by (2). We show that $E$ is the offset adjacency matrix of some totally ranked tree shape, i.e., $E_{i,j}=A_{j,i+1}$ where $A$ is the adjacency matrix. Recall that in the adjacency matrix, entry $A_{j,i+1}=1$ if nodes with ranks $j$ and $i+1$ are connected by an edge. This requires verifying the following conditions for $E$:
@@
We can determine all F-matrices by filling the remaining entries in order of $F_{2,0}$, $F_{3,0}$, and $F_{3,1}$. For $F_{2,0}$ we have two options: 0 or 1. Suppose we select $F_{2,0}=0$. Moving to $F_{3,0}$, we are forced to select $F_{3,0}=0$ by Constraint 2. Lastly, for $F_{3,1}$ our options are 1 or 2, both of which yield valid F-matrices. One can verify that if we instead begin with $F_{2,0}=1$, the remaining entries work out in a similar fashion.
@@
For $F_{2,0}$ we have two options, 0 or 1; suppose we take $F_{2,0}=0$. We are forced to have $F_{2,1}=2$ by Condition 33(b). Next we must take $F_{2,2}=2$, as $F_{2,2}=4$ yields the contradiction $3=F_{3,2}\leq F_{3,3}=1$ by Conditions 3(a,b). Additionally, we are forced to take $F_{3,0}=0$ by Constraint 2. In
@@
This example shows how the strategy of filling rows one by one from top to bottom (i.e., traversing entries $F_{0,0},F_{1,0},F_{1,1},F_{2,0},\cdots,F_{2n-3,2n-4},F_{2n-3,2n-3}$) produces all F-matrices, in both the isochronous and heterochronous case, but additional constraints are necessary to prevent invalid F-matrices in the heterochronous case. Specifically, some combinations of values for $F_{i,j}$ and $F_{i,i-1}$ from Conditions 33(c) and 33(b) in Theorem 2 may conflict with item 1. In the example, the invalid combination is $F_{i,j}=F_{3,1}=2$ and $F_{i,i-1}=F_{3,2}=1$.
For the remainder of this section, we describe a matrix-filling strategy that does not lead to contradictions in the heterochronous case. As the subdiagonal entries are determined by the diagonal entries, we first verify that any choice of diagonal entries by Condition 33(a) and an additional constraint yields at least one valid $\mathbf{F}$-matrix. That is to say, when solving the system of inequalities in Theorem 2, we may select values for the diagonal without backtracking.
@@
The following lemma is necessary for extending an F-sequence by one entry.
Lemma 1.
If $f_{i,j}$ is an $(n,N,M)$ F-sequence with M<N2M<N-2, then
$f_{N,M}\leq f_{N-1,M}\leq f_{N-1,M+1}.$
Proof.
By Definition 2.2, $f_{N,M}\leq U_{f}(N,M)\leq f_{N-1,M}$. We handle the remaining bound in three cases. First note that by Definitions 2.3 and 2.1,
$f_{N-2,N-3}=f_{N-3,N-3}-1=(f_{N-2,N-2}\pm 1)-1.$
When M=N3M=N-3 and $f_{N-2,N-3}=f_{N-2,N-2}$, by Definition 2.3,
$f_{N-1,M}=f_{N-1,N-3}=f_{N-2,N-2}-1=f_{N-1,N-2}=f_{N-1,M+1}.$
When M=N3M=N-3 and $f_{N-2,N-3}=f_{N-2,N-2}-2$, Definitions 2.2 and 2.3 give
$<f_{N-1,M+1}.$
Lastly, when M<N3M<N-3, by Definition 2.2,
$f_{N-1,M+1}\geq L_{f}(N-1,M+1)\geq f_{N-1,M}.$
We show how to extend an $(n,N,M)$ F-sequence to a $(2n-2)\times(2n-2)$ F-matrix and that every F-matrix appears in such a way.